OpenStudy (anonymous):

I am having trouble writing an equation for a word problem. The problem says that kay wants to rent a car. She has two plans to choose from. plan one has a fee of $65. and 40 cents for additional miles. Plan two has no fee but costs 60 cents a mile. how many miles will she drive in order for the plans to cost the same? 65 + 40x=60x ???????

OpenStudy (anonymous):

65 = 60x - 40x 65 = 20x 65/20 = x 3.25 = x

OpenStudy (anonymous):

This is wrong......... Units should be same.....

OpenStudy (anonymous):

65 + 0.40x = 0.60x 65 = 0.60x - 0.40x 65 = 0.20x 65/0.20 = x 325 = x she will have to drive 325 km for both amounts to be equal..